# Exploring Ordinal Bias in Action Recognition for Instructional Videos

### v1

@@ 1 Introduction
Action recognition in instructional videos has witnessed remarkable progress, primarily driven by models that excel in curated benchmark datasets (Farha & Gall, 2019; Ishikawa et al., 2021; Li et al., 2020; Yi et al., 2021). However, these datasets often present a limited view of real-world variability by favoring specific, repeated action sequences. In contrast, real-world instructional videos, such as those for cooking or physical exercise, frequently present more diverse action sequences than those captured in benchmark datasets. As a result, models trained on conventional benchmarks tend to exploit these spurious ordinal patterns that we refer to as ordinal bias.
We observe that existing datasets (Fathi et al., 2011; Stein & McKenna, 2013; Kuehne et al., 2014) demonstrate biased action sequences, which lead the model to suffer from spurious correlations. As shown in Figure 1, the dataset exhibits a dominant occurrence of the action ‘Take’ followed by ‘Background.’ This biases the model toward learning spurious correlations, causing it to predict ‘Background’ as the next action rather than relying on visual inputs to correctly predict ‘Open.’ This raises concerns about the reliability of evaluations and the risk of overestimating the performance of the action recognition models.
@@ 2.1 Recognition of Actions in Instructional Videos
Instructional video analysis has emerged as a prominent area of research in the field of video comprehension. In particular, multiple datasets of instructional videos (Fathi et al., 2011; Stein & McKenna, 2013; Kuehne et al., 2014) have been introduced, offering extensive contextual information on human activities. Although various methods have been proposed (Farha & Gall, 2019; Li et al., 2020; Ishikawa et al., 2021; Yi et al., 2021; Liu et al., 2023), and Li et al. (2022) explores the understanding of ordinal action by explicitly modeling inter-action context, the ordinal bias issue can overestimate their performance. This comes from their exploitation of the action sequence observed during training stage.
@@ 2.2 Bias in Action Recognition
Recent work has advanced reliable video systems and evaluations across various video understanding tasks. Otani et al. (2020); Yuan et al. (2021); Jung et al. (2024) have analyzed the reliable evaluation of temporal understanding, and Li et al. (2018); Hara et al. (2021) have raised the bias problem in action recognition. Consequently, a line of works (Zhai et al., 2023; Li et al., 2024) have explored the dual challenges of background and foreground biases, demonstrating that action recognition models can be inadvertently biased by static and dynamic cues. Although the previous works (Nam et al., 2020; Duan et al., 2022) have addressed the bias by retraining the model, these approaches are computationally expensive and overlook the underlying imbalance of the data set. In contrast, we directly address the issue by manipulating the video data itself and provide insights into how current models perform if they are trained with these video variants.
@@ 3.1 Dataset
We utilize three action recognition datasets: Georgia Tech Egocentric Activities (GTEA) (Fathi et al., 2011), 50Salads (Stein & McKenna, 2013), and Breakfast (Kuehne et al., 2014). GTEA includes 28 videos depicting daily kitchen activities, featuring 11 action categories. Each video has an average of 20 action units and a duration of approximately 30 seconds. 50Salads contains 50 videos of actors preparing salads in various kitchen environments, with more than 20 actors participating. The videos in 50Salads are more than six minutes long and cover 17 action categories. Breakfast consists of over 1700 videos that contain breakfast preparation scenes and has 48 action categories. This dataset has the most complex labeling scheme among the three datasets.
@@ 3.2 Action Recognition Models
We consider five distinct models: MS-TCN (Farha & Gall, 2019), MS-TCN++ (Li et al., 2020), ASRF (Ishikawa et al., 2021), and DiffAct (Liu et al., 2023). MS-TCN employs a multi‑stage architecture with dilated temporal convolutions and a smoothing loss to iteratively refine frame‑level predictions. MS-TCN++ extends this approach by integrating dual dilated layers that capture both local and global contexts while decoupling prediction generation from refinement. ASRF improves segmentation quality by adding an auxiliary branch that explicitly regresses action boundaries to mitigate over‑segmentation errors. ASFormer leverages a Transformer‑based framework augmented with temporal convolutions and a hierarchical representation pattern for iterative prediction refinement. Lastly, DiffAct formulates action segmentation as a conditional sequence generation task that iteratively denoises a noisy action sequence by leveraging priors such as positional, boundary, and relational cues. For our experiments, we utilized I3D (Carreira & Zisserman, 2017) video features, pre-trained on the Kinetics dataset (Kay et al., 2017), and a single NVIDIA RTX 3090.
@@ 5.2 Impact of Additional Training
Table 2: Accuracy of the models trained with additional datasets. O/O: Trained and tested on the original dataset; C/O: Trained on the combined dataset, tested on the original dataset; C/S: Trained on the combined, tested on the sequence shuffling; C/M: Trained on the combined, tested on the action masking.
In many bias-related problems, the incorporation of additional data helps alleviate bias. Therefore, we investigate whether training models with an additional augmented dataset can mitigate ordinal bias in action recognition. To this end, we designed a curriculum learning-like strategy by sequentially training the model on three variants of the dataset: the original, the masked, and the shuffled versions. This progression is intended to gradually expose the model to increasing difficulty levels and reduce its reliance on spurious correlations. However, as highlighted in Section 5.1, the use of action masking is not suitable for model training. To overcome this, we used a strategy for action masking inspired by Masked Language Modeling (Devlin, 2018), where actions are randomly hidden with a likelihood of 15%, instead of our original approach. In contrast, for sequence shuffling, we utilized the method we had proposed. Consequently, every augmented dataset retains its original size, allowing the model to be trained on a dataset that is triple the size of the initial one, which we refer to as the ‘Combined’ dataset.

### v2

@@ 1 Introduction
Action recognition in instructional videos has witnessed remarkable progress, primarily driven by models that excel in curated benchmark datasets (farha2019ms; ishikawa2021alleviating; li2020ms; yi2021asformer). However, these datasets often present a limited view of real-world variability by favoring specific, repeated action sequences. In contrast, real-world instructional videos, such as those for cooking or physical exercise, frequently present more diverse action sequences than those captured in benchmark datasets. As a result, models trained on conventional benchmarks tend to exploit these spurious ordinal patterns that we refer to as ordinal bias.
We observe that existing datasets (fathi2011learning; stein2013combining; kuehne2014language) demonstrate biased action sequences, which lead the model to suffer from spurious correlations. As shown in Figure 1, the dataset exhibits a dominant occurrence of the action ‘Take’ followed by ‘Background.’ This biases the model toward learning spurious correlations, causing it to predict ‘Background’ as the next action rather than relying on visual inputs to correctly predict ‘Open.’ This raises concerns about the reliability of evaluations and the risk of overestimating the performance of the action recognition models.
@@ 2.1 Recognition of Actions in Instructional Videos
Instructional video analysis has emerged as a prominent area of research in the field of video comprehension. In particular, multiple datasets of instructional videos (fathi2011learning; stein2013combining; kuehne2014language) have been introduced, offering extensive contextual information on human activities. Although various methods have been proposed (farha2019ms; li2020ms; ishikawa2021alleviating; yi2021asformer; liu2023diffusion), and li2022bridge explores the understanding of ordinal action by explicitly modeling inter-action context, the ordinal bias issue can overestimate their performance. This comes from their exploitation of the action sequence observed during training stage.
@@ 2.2 Bias in Action Recognition
Recent work has advanced reliable video systems and evaluations across various video understanding tasks. otani2020uncovering; yuan2021closer; jung2024consistency have analyzed the reliable evaluation of temporal understanding, and li2018resound; hara2021rethinking have raised the bias problem in action recognition. Consequently, a line of works (soar2023; li2024fairaction) have explored the dual challenges of background and foreground biases, demonstrating that action recognition models can be inadvertently biased by static and dynamic cues. Although the previous works (nam2020learning; duan2022mitigating) have addressed the bias by retraining the model, these approaches are computationally expensive and overlook the underlying imbalance of the data set. In contrast, we directly address the issue by manipulating the video data itself and provide insights into how current models perform if they are trained with these video variants.
@@ 3.1 Dataset
We utilize three action recognition datasets: Georgia Tech Egocentric Activities (GTEA) (fathi2011learning), 50Salads (stein2013combining), and Breakfast (kuehne2014language). GTEA includes 28 videos depicting daily kitchen activities, featuring 11 action categories. Each video has an average of 20 action units and a duration of approximately 30 seconds. 50Salads contains 50 videos of actors preparing salads in various kitchen environments, with more than 20 actors participating. The videos in 50Salads are more than six minutes long and cover 17 action categories. Breakfast consists of over 1700 videos that contain breakfast preparation scenes and has 48 action categories. This dataset has the most complex labeling scheme among the three datasets.
@@ 3.2 Action Recognition Models
We consider five distinct models: MS-TCN (farha2019ms), MS-TCN++ (li2020ms), ASRF (ishikawa2021alleviating), and DiffAct (liu2023diffusion). MS-TCN employs a multi‑stage architecture with dilated temporal convolutions and a smoothing loss to iteratively refine frame‑level predictions. MS-TCN++ extends this approach by integrating dual dilated layers that capture both local and global contexts while decoupling prediction generation from refinement. ASRF improves segmentation quality by adding an auxiliary branch that explicitly regresses action boundaries to mitigate over‑segmentation errors. ASFormer leverages a Transformer‑based framework augmented with temporal convolutions and a hierarchical representation pattern for iterative prediction refinement. Lastly, DiffAct formulates action segmentation as a conditional sequence generation task that iteratively denoises a noisy action sequence by leveraging priors such as positional, boundary, and relational cues. For our experiments, we utilized I3D (carreira2017quo) video features, pre-trained on the Kinetics dataset (kay2017kinetics), and a single NVIDIA RTX 3090.
@@ 5.2 Impact of Additional Training
Table 2: Accuracy of the models trained with additional datasets. O/O: Trained and tested on the original dataset; C/O: Trained on the combined dataset, tested on the original; C/S: Trained on the combined, tested on the sequence shuffling dataset; C/M: Trained on the combined, tested on the action masking dataset.
In many bias-related problems, the incorporation of additional data helps alleviate bias. Therefore, we investigate whether training models with an additional augmented dataset can mitigate ordinal bias in action recognition. To this end, we designed a curriculum learning-like strategy by sequentially training the model on three variants of the dataset: the original, the masked, and the shuffled versions. This progression is intended to gradually expose the model to increasing difficulty levels and reduce its reliance on spurious correlations. However, as highlighted in Section 5.1, the use of action masking is not suitable for model training. To overcome this, we used a strategy for action masking inspired by Masked Language Modeling (devlin2018bert), where actions are randomly hidden with a likelihood of 15%, instead of our original approach. In contrast, for sequence shuffling, we utilized the method we had proposed. Consequently, every augmented dataset retains its original size, allowing the model to be trained on a dataset that is triple the size of the initial one, which we refer to as the ‘Combined’ dataset.